From nch@roe.ac.uk Sun Feb 12 15:08:34 2006 Date: Fri, 10 Feb 2006 16:29:10 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , John Taylor , Jim Emerson , Malcolm Stewart , Martin Hill , Mike Irwin , Peredur Williams , Stephen Warren Subject: WFAU WSA weekly meeting minutes, 10th Feb 2006 Minutes of WFCAM Science Archive meeting: 10th February 2006 ------------------------------------------------------------- ------------------------------------------------------------- Present: NCH, ETWS, RSC, PMW, JB, RGM, MAR Apologies: JPE, JMS, AL, JDT, NJC DONM: 10am, Friday 17th February 2006, plate library Actions discharged: ------------------- ACTION: ETWS to nudge U.Chicago over SDSS-DR3 again. Discharged; no word as yet. Incidentally, an email from Jim Gray and the JHU mob notes that BestDR4 is 1.7TB in size. ACTION: PMW to chase JMS/JPE for VDFS UK review TOR Discharged: JMS reminded PMW and NCH that this probably originated in draft form as a result of the last VDUC meeting. Actions partly discharged but continuing: ----------------------------------------- ACTION: JDT to set up a Datascope Launcher for the WSA on the AG workbench pages. Continues. Actions carried forward from 27/01/06 meeting: ---------------------------------------------- ACTION: NCH to investigate the usefulness of a NAS solution for medium term archive mass storage. - CONTINUES Specific points and new actions: -------------------------------- Project management: JMS noted the ongoing need to sort out the UK review. Organising the next VDMT is proving to be problematic... WFCAM update: See the stop press pages linked from the TWiki for latest news. Comments and issues arising from CASU fortnightly minutes: No new minutes as of 10/2/06 AM. Networking: Nothing new this week. WSA Operations: PMW noted that the premises and IT support teams will be doing electrical saftey checking in C1 over the weekend of 25/26th Feb. ACTION: MAR to put a note on the "downtime" web page to inform users of the small interuption to normal service. JB reported: "Reran CU7 for the DXS, CU16 for the UDS and DXS, running CU18, copying EDRplus and EDR to Amenhotep and made tape backup copies then attached them to the SQL Server. Finished catching up the system tape backups and have also (in addition to the EDR tape backups) done GLIMPSE and its flat flies, some more of the old log files and have a plan for WORLDR2, UKIDSSR1 and UKIDDSR2. Helped MAR discover the reason for Amenhotep being slow (user query that was mistyped and badly contructed). Replaced and rebuilt a broken disc on Ahmose." Hardware: PMW noted that the new filestore server djedefre will be delivered next week. Software: RSC reported: "Improved the efficiency of the script that creates the Provenance table, with the aid of some complex SQL courtesy of NCH. Determined that the extent of the incomplete provenance information in the database is limited to just the partially transferred 05B data and the early 05A data, which has always been known to be incomplete. Continued testing of the new Cu16 script, which now appears to be working fine for all databases, other than Sloan DR2, which requires additional attributes in the neighbour table. I'm in the process of investigating the most sensible way of handling this issue." ETWS reported: "Had a look through all the browser pages and schemas to fix typos, inconsistencies and duplicate entries. Replaced compressed images (jpegs) of deprecated detector frames with a default image after Johann and IT sorted out the disk space problem on disk06 caused by Nick. Started to have a look into the parser to reflect different versions of browser pages for different releases." NCH has been finalising the preparation of the UKIDSS EDR databases for release today. Survey Data Release: EDR release occured this day, the 10th February 2006. The team discussed the issue of the constantly changing DB schema and how to document this, at the same time minimising confusion for users. After an exchange of ideas, the decision was made to put an intervening page between the browser link and the actual browser to direct users to the right one, with an obvious default to the current release browser for folks who don't know any better. ACTION: ETWS and MAR to enhance the schema browser pages to cope with multiple browser versions, supporting previous releases. NCH noted that a number of issues and bugs had cropped up recently, and that there must be a clear note of these on the website. ACTION: NCH to put a list of known issues and bugs up under the archive release history. Non-Survey Data Release: MAR noted two new 05A non-survey registrations; NCH suggested that JB prepare their release DBs as requested. Astrogrid deployment: Nothing new this week. Miscellaneous: Nothing else to report this week.